### v1

# Batch-of-Thought: Cross-Instance Learning for Enhanced LLM Reasoning

Abstract
Current Large Language Model reasoning systems process queries independently, discarding valuable cross-instance signals such as shared reasoning patterns and consistency constraints. We introduce Batch-of-Thought (BoT), a training-free method that processes related queries jointly to enable cross-instance learning. By performing comparative analysis across batches, BoT identifies high-quality reasoning templates, detects errors through consistency checks, and amortizes computational costs. We instantiate BoT within a multi-agent reflection architecture (BoT-R), where a Reflector performs joint evaluation to unlock mutual information gain unavailable in isolated processing. Experiments across three model families and six benchmarks demonstrate that BoT-R consistently improves accuracy and confidence calibration while reducing inference costs by up to 61%. Our theoretical and experimental analysis reveals when and why batch-aware reasoning benefits LLM systems.
Batch-of-Thought: Cross-Instance Learning for Enhanced LLM Reasoning
Xuan Yang1,2, Furong Jia1, Roy Xie1, Xiong Xi2, Hengwei Bian2, Jian Li2, Monica Agrawal1 1Duke University 2ByteDance Inc. {xuan.yang, flora.jia, ruoyu.xie, monica.agrawal}@duke.edu {xi.xiong, hengwei.bian, limingjun.tsinghua}@bytedance.com
1 Introduction
Large Language Models (LLMs) openai2023gpt4; madaan2023selfrefine; wei2022cot; shinn2023reflexion; yao2023react have achieved strong performance across diverse tasks and are increasingly applied in domains such as medical reasoning, question answering, and scientific problem solving (Singhal2025TowardEMA; mcduff2025towards; nori2025sequential; haas2025simpleqa; wang2023scibench; sun2024scieval). However, producing reliable answers with well-calibrated confidence remains a challenge  xiong2023uncertainty; ji2023hallucination_survey. LLMs often assign high confidence to incorrect answers, which undermines their practical deployment in high-stakes applications where accuracy and reliable uncertainty quantification are essential.
Figure 1: Batch-of-Thought reflection framework. An Actor generates initial responses for a batch of queries. A Reflector then jointly evaluates all responses through comparative analysis, determining whether each should be finalized or refined with feedbacks.
Multi-agent LLM systems (li2023camel; chan2023chateval; guo2024large) extend single-model capabilities through specialized roles and iterative refinement. Despite architectural diversity, existing approaches share a fundamental limitation: they process queries independently. While computationally straightforward, this paradigm discards valuable cross-instance signals. When queries share domain characteristics or structural patterns, isolated processing prevents the system from identifying outliers through comparative assessment, propagating validated knowledge from confident instances to uncertain ones, or detecting errors that emerge only through cross-instance consistency checks.
We introduce Batch-of-Thought (BoT), a training-free framework that processes related queries jointly to enable cross-instance learning and comparative reasoning. Our key insight is that batch-level reasoning unlocks mutual information gain unavailable in isolated processing. By treating queries as a cohort rather than independent instances, BoT enables comparative analysis, reasoning pattern identification, and distributional uncertainty calibration.
To illustrate this principle, consider fraud detection: evaluating a single seller in isolation provides limited signal. Examining a cohort simultaneously reveals recurring suspicious patterns, distinguishes legitimate domain practices from genuine anomalies, and enables comparative evidence assessment. This principle mirrors James–Stein estimation james1961quadratic; stein1956inadmissibility; efron1977steinsparadox: pooling information across similar instances improves individual estimates through shrinkage toward the cohort distribution (see Appendix C for theoretical analysis).
We instantiate BoT within a multi-agent reflection architecture madaan2023selfrefine; shinn2023reflexion, termed BoT-R, though the principle generalizes to other frameworks including Plan-and-Act Erdogan2025PlanAndAct and Multi-Agent Debate Du2023MultiAgentDebate; Liang2023MAD. In BoT-R, an Actor generates answer-rationale pairs for a batch of queries, then a Reflector performs joint evaluation through comparative analysis—identifying inconsistencies, extracting shared domain knowledge, assessing relative quality, and suggesting refinements. This approach simultaneously improves reasoning quality and computational efficiency by amortizing reflection overhead across the batch. We summarize our main contributions as follows:
  1. 1.
    We propose Batch-of-Thought (BoT), a training-free method that enhances LLM reasoning by processing related queries as cohesive batches, enabling cross-instance learning unavailable in isolated processing.
  2. 2.
    We instantiate BoT in a reflection-based multi-agent system and conduct experiments across six benchmarks and three model families, demonstrating consistent accuracy improvements and 46.9% average cost reduction.
  3. 3.
    We theoretical and experimental analyze how task characteristics and batch composition influence BoT’s effectiveness, revealing that interpretive domains benefit substantially from comparative reasoning while symbolic tasks require careful batch design.
  4. 4.
    We introduce the Seller Fraud Detection benchmark for evaluating agentic reasoning in high-stakes scenarios, which we will publicly release.
Algorithm 1 BoT-R
Batch $B=\{x_{i}\}_{i=1}^{N}$, Actor $\mathcal{A}$, Reflector $\mathcal{R}$, tool set $\mathcal{T}$, max outer rounds $T$, max tool calls $K$
Final answers $\{a_{i}\}_{i=1}^{N}$ with confidences $\{u_{i}\}_{i=1}^{N}$
Initialize $c_{i}\leftarrow\varnothing$, $u_{i}\leftarrow 0$ for all $i\in[N]$; active set $S\leftarrow[N]$
for $t=1$ to $T$ do
  (Parallel) $(a_{i},\rho_{i},\mathsf{traj}_{i})\leftarrow\mathcal{A}(x_{i},\mathcal{T},c_{i},K)$ for all $i\in S$
  Build reflective context $\mathcal{C}\leftarrow\langle(x_{i},a_{i},\rho_{i},\mathsf{traj}_{i})\rangle_{i=1}^{N}$
  (Joint) $(u_{i},r_{i},c_{i})\leftarrow\mathcal{R}(\mathcal{C},i)$ for all $i\in[N]$
  if $\forall i:\ r_{i}=0$ then break
  $S\leftarrow\{\,i\mid r_{i}=1\}$
return $\{a_{i}\},\{u_{i}\}$
Model Method FraudDet GPQA Winogrande MedQA PubMedQA SMS Spam
$n=1793$ $n=448$ $n=1267$ $n=1273$ $n=1000$ $n=1510$
GPT-4o ReAct 0.685 0.439 0.872 0.878 0.679 0.796
Reflection 0.693 0.459 0.879 0.901 0.667 0.854
BOT-R 0.740 0.488 0.890 0.904 0.698 0.887
Llama-3.3-70B ReAct 0.635 0.494 0.831 0.783 0.753 0.920
Reflection 0.679 0.504 0.853 0.797 0.755 0.925
BOT-R 0.713 0.516 0.862 0.804 0.757 0.923
Qwen3-Next-80B ReAct 0.633 0.560 0.823 0.814 0.732 0.946
Reflection 0.639 0.636 0.869 0.846 0.681 0.900
BOT-R 0.660 0.657 0.874 0.860 0.704 0.919
Table 1: Performance comparison of reasoning methods across various base models and datasets (number of queries listed as $n$). Scores represent accuracy. The best result in each setting is highlighted in bold.
2 Methods
Batch-of-Thought (BoT) is a training-free, model-agnostic method that jointly processes batches of queries to improve reasoning quality, confidence calibration, and computational efficiency. We formalize the approach and describe its instantiation within a multi-agent reflection architecture.
2.1 Problem Formulation
Let $\mathcal{X}$ and $\mathcal{Y}$ denote input and output spaces. Queries arrive in batches $B=\{x_{i}\}_{i=1}^{N}\subset\mathcal{X}$. We employ a two-agent architecture with iterative refinement:
Actor $\mathcal{A}$.
A ReAct agent yao2023react that interleaves reasoning traces with tool execution to generate answer-rationale pairs. At iteration $t$, given query $x_{i}$ and optional critique $c_{i}^{(t-1)}$ from the previous round:
$(a_{i}^{(t)},\rho_{i}^{(t)})=\mathcal{A}(x_{i},c_{i}^{(t-1)};\text{tools}),\quad a_{i}^{(t)}\in\mathcal{Y},$ (1)
where $c_{i}^{(0)}=\varnothing$ for the initial iteration.
Reflector $\mathcal{R}$.
A reflection agent madaan2023selfrefine; shinn2023reflexion that evaluates context $\mathcal{C}^{(t)}$ containing all current answer-rationale pairs. For each query $i$, it produces:
$(r_{i}^{(t)},u_{i}^{(t)},c_{i}^{(t)})=\mathcal{R}(\mathcal{C}^{(t)},i),$ (2)
where $r_{i}^{(t)}\in\{0,1\}$ indicates whether query $i$ requires refinement, $u_{i}^{(t)}\in[0,1]$ is a confidence score, and $c_{i}^{(t)}$ is an actionable critique. If $r_{i}^{(t)}=1$, the query proceeds to iteration $t+1$ with critique $c_{i}^{(t)}$; otherwise, $a_{i}^{(t)}$ is finalized.
Objective.
Improve (i) task accuracy, (ii) confidence calibration, and (iii) computational efficiency
2.2 Batch-of-Thought (BoT)
Formalization.
Standard per-instance reflection constructs $N$ independent contexts:
$\mathcal{C}_{i}^{\text{ind}}=\langle(x_{i},a_{i},\rho_{i})\rangle,$ (3)
evaluating each query in isolation. BoT instead constructs a single shared context:
$\mathcal{C}^{\text{BoT}}=\langle(x_{1},a_{1},\rho_{1}),\ldots,(x_{N},a_{N},\rho_{N})\rangle,$ (4)
and performs joint evaluation:
$\bigl\{(r_{i},u_{i},c_{i})\bigr\}_{i=1}^{N}=\mathcal{R}\bigl(\mathcal{C}^{\text{BoT}}\bigr).$ (5)
Cross-instance mechanisms.
The shared context enables three synergistic mechanisms: (1) Outlier detection: $\mathcal{R}$ identifies answers that appear plausible in isolation but are inconsistent with peer patterns, propagating high-quality reasoning templates via critiques $\{c_{i}\}$. (2) Distributional calibration: Confidence scores $u_{i}$ are calibrated relative to batch statistics $\phi(\mathcal{C}^{\text{BoT}})$ rather than assessed independently, improving uncertainty quantification. (3) Computational amortization: Evaluation rubrics are encoded once per batch, reducing input costs, and joint evaluation enables more accurate refinement decisions, reducing unnecessary Actor-Reflector loops.
The complete BoT-R workflow is detailed in Algorithm 1, which alternates between Actor generation and Reflector evaluation steps until convergence or maximum iterations.
Theoretical foundation.
Appendix C establishes formal guarantees through information-theoretic and statistical analysis, demonstrating that BoT achieves a Pareto improvement over independent processing: simultaneously enhancing accuracy and reducing computational cost.
3 Experiments
3.1 Experimental Setup
We evaluate BoT on six datasets, including five public benchmarks and one newly curated corpus, using both API-based and open-source large language models. Full experimental details are provided in Appendix A.
Datasets.
Our evaluation covers diverse reasoning and decision-making tasks: GPQA (rein2024gpqa), WinoGrande–debiased (sakaguchi2021winogrande), PubMedQA (jin2019pubmedqa), MedQA (USMLE) (jin2021disease), MMLU (hendrycks2020measuring), SMS Spam Detection (yang2024ad), and a newly curated dataset fraud-seller detection dataset (Appendix G). Together, these benchmarks span scientific reasoning, commonsense inference, biomedical QA, broad academic knowledge, and real-world anomaly detection.
Metrics.
We evaluate (i) task accuracy, (ii) token efficiency (input token count, output token count, and total), and (iii) confidence calibration using two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation between the confidence distributions of correct vs. incorrect predictions, and (b) Expected Calibration Error (ECE; guo2017calibration).
Baselines.
To isolate BoT’s contribution, we compare against two training-free reasoning baselines: ReAct (yao2023react), which performs standard single-instance reasoning with optional tool augmentation, and Reflection (shinn2023reflexion; madaan2023selfrefine), a multi-agent framework employing per-instance self-critique and revision. We then augment the same Actor with BoT’s joint, batch-aware reflection to obtain BoT-R. All other factors remain constant across methods.
Models.
We report results from both API and open-source models: GPT-4o-2024-11-20 hurst2024gpt, Llama-3.3-70B llama3 and Qwen3-Next-80B yang2025qwen3. For the Fraud Detection dataset, we enable Brave Search and the Brave Summarizer as external tools for retrieval and grounding brave_api_docs
SMS_Spam GPQA Winogrande
Method Cost $\Delta\%$ Cost $\Delta\%$ Cost $\Delta\%$
Actor
Reflection \$3.61 \$4.70 \$2.57
BOT (4) \$2.41 33.25% \$3.68 21.71% \$2.27 11.75%
BOT (8) \$2.06 42.96% \$3.42 27.24% \$1.99 22.54%
Reflector
Reflection \$6.39 \$2.76 \$3.40
BOT (4) \$2.44 61.85% \$1.49 45.99% \$1.97 42.10%
BOT (8) \$1.84 71.12% \$1.17 57.65% \$1.52 55.31%
Total
Reflection \$10.00 \$7.46 \$5.97
BOT (4) \$4.85 51.52% \$5.17 30.68% \$4.24 29.04%
BOT (8) \$3.9 60.95% \$4.59 38.48% \$3.51 41.21%
Table 2: Total token cost and relative reduction ($\Delta\%$) for each dataset(GPT-4o). $\Delta\%$ is computed for BOT(4) and BOT(8) relative to Reflection cost.
Method SMS_Spam GPQA Winogrande
KS↑ ECE↓ KS↑ ECE↓ KS↑ ECE↓
ReAct 0.256 0.176 0.181 0.372 0.273 0.113
Reflect 0.360 0.104 0.265 0.329 0.376 0.035
BoT-R 0.633 0.063 0.368 0.317 0.442 0.013
Table 3: Confidence calibration across datasets. Each entry reports the KS statistic between the confidence distributions of correct vs. incorrect answers and the ECE score. Higher KS (↑) is better, while lower ECE (↓) is better.
Method med&bio hum social math sci
ReAct 0.887 0.805 0.915 0.763 0.797
Reflection 0.886 0.825 0.915 0.865 0.843
BoT-R 0.891 0.837 0.922 0.853 0.832
Table 4: MMLU dataset subject-wise accuracy (GPT-4o). The highest score for each subject is in bold.
3.2 Main Results
As shown in Table 1, we compare task performance after integrating the proposed BoT method into the reflection framework. Across three backbones, BoT-R is consistently competitive and typically the strongest overall variant, improving over both ReAct and standard Reflection on most dataset–model pairs. The gains are most visible on higher-variance, decision-heavy tasks where per-instance reflection can be brittle. For example, under GPT-4o, BoT-R improves FraudDet and GPQA by +4.7+4.7 and +2.9+2.9 accuracy points over Reflection, respectively, and yields an average improvement of +2.6+2.6 points across all six datasets. In contrast, on near-saturated benchmarks where base performance is already high, the accuracy headroom is limited, and the improvements are naturally smaller, suggesting that BoT is most beneficial when cross-instance comparison provides additional corrective signal.
We further evaluate efficiency using a unified reference pricing scheme based on production-grade GPT-4o pricing. Table 2 shows that BoT-R substantially reduces overall token cost, achieving 46.9% average savings across the three representative benchmarks at batch size $8$, and up to 61% reduction on SMS_Spam. These reductions indicate that batch-aware reflection effectively amortizes reflective reasoning across instances while improving task performance.
Finally, Table 3 shows that BoT-R improves confidence reliability under GPT-4o. It increases KS and reduces ECE across all three datasets (e.g., on SMS_Spam, KS $0.360\rightarrow 0.633$ and ECE $0.104\rightarrow 0.063$). This is consistent with the collective-signal perspective: when the effective sample size $N_{\mathrm{eff}}$ is meaningfully above $1$ (as expected for moderate correlation at $N\in\{4,8\}$), batch-level consensus provides a stronger signal for separating correct from incorrect predictions, which directly predicts higher KS and improved calibration.
Overall, the results support a clear conclusion, consistent with our theoretical analysis: batch-aware reflection yields a favorable accuracy–cost–calibration trade-off, with robust accuracy gains and calibration improvements on diverse tasks and consistent efficiency. Additional details are provided in AppendixB̃.
4 Discussions
RQ1: Which domains benefit most from BoT?
Table 4 shows that BoT-R yields the largest gains on interpretive and judgment-driven domains, including humanities, social sciences, and medicine & biology. These tasks admit multiple plausible reasoning paths and partial cues, making comparative evaluation across instances especially informative. In contrast, domains dominated by exact symbolic derivation, such as mathematics and parts of the physical sciences, exhibit marginal or slightly negative changes. This suggests that batch-level consensus is less effective when correctness depends on exact derivation rather than comparative plausibility.
From a theoretical perspective, this pattern aligns with the coherence and informativeness conditions in Section C. Interpretive tasks tend to satisfy moderate similarity and moderate error correlation, yielding a meaningful effective sample size and enabling collective signal amplification. By contrast, symbolic domains often violate these assumptions: small derivation errors are highly correlated within a batch, limiting the benefit of aggregation and occasionally amplifying shared mistakes.
RQ2: How does batching strategy influence performance?
Our results indicate that how queries are batched matters, but less than might be expected. Sequential batching—grouping queries in their natural order—already delivers consistent improvements over instance-wise reflection across all six datasets (Appendix E). This suggests that many benchmarks exhibit latent topical or structural coherence, allowing BoT-R to extract useful cross-instance signals even without explicit clustering.
Semantic batching further improves performance on heterogeneous datasets such as FraudDet and Winogrande. By increasing within-batch similarity, embedding-based grouping reduces noise in comparative evaluation and strengthens cross-instance signals. These gains follow our theoretical prediction that coherent batches yield stronger cross-instance signals.
Batch size introduces an additional trade-off. While larger batches theoretically provide richer comparative context, empirical results show a non-monotonic relationship between batch size and performance. Moderate batch sizes ($N\in\{4,8\}$) offer the best accuracy–efficiency balance. Larger batches are constrained by (i) context window saturation, which forces rationale compression and degrades fine-grained reasoning, and (ii) increased heterogeneity, which dilutes informative cross-instance comparisons. As a result, performance often peaks at intermediate batch sizes despite stronger theoretical aggregation benefits.
Implications and outlook.
Overall, these findings suggest that BoT is most effective when batches exhibit sufficient, but not excessive, coherence, and when tasks benefit from comparative judgment rather than exact symbolic derivation. Importantly, the robustness of sequential batching indicates that BoT-R remains practical in streaming or latency-sensitive settings, where semantic clustering may be infeasible. Future work may explore adaptive cohorting strategies that dynamically balance coherence, batch size, and latency, as well as extensions to domains requiring symbolic guarantees.
5 Conclusion
We introduced Batch-of-Thought (BoT), a training-free, model-agnostic approach that processes related queries as a batch so an agent can perform comparative analysis, share knowledge across items, and produce better-calibrated confidence while amortizing computation. BoT yields higher accuracy, lower token cost, and improved calibration across settings, including our proposed Seller Fraud Detection benchmark.
Limitations
The efficacy of Batch-of-Thought (BoT) is subject to several constraints. First, performance depends on batch formation: the core comparative reasoning assumes within-batch semantic relatedness. Poorly formed cohorts can cause negative transfer, degrading both calibration and accuracy. Second, BoT inherits the base model’s context limits. For long-context queries, concatenating multiple items may approach or exceed the window, leading to truncation or failures. Finally, while we instantiate BoT in an Actor–Reflector architecture and the idea generalizes naturally to other multi-agent designs (e.g., Plan-and-Act, Debate), empirical validation of such integrations remains open. Systematically assessing the portability of batch-aware reasoning across alternative collaborative frameworks is an important direction for future work.
Ethics Statement
In writing this paper, we used an AI assistant to correct grammatical errors. During the coding process, we utilized AI tools for code completion. All datasets and models used in our experiments are publicly accessible. Our newly released Seller Fraud Detection benchmark contains only publicly available information and does not include any private or sensitive data. The Seller Fraud Detection benchmark was developed with human expert annotation. All annotators were compensated fairly for their time and expertise at rates exceeding standard professional compensation in their region. Annotators were provided with clear guidelines and had the option to decline participation at any time.
Appendix A Experiment Settings
Confidence Calibration Metrics.
We use two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation, which measures the maximum difference between the cumulative distributions of confidence scores for correct versus incorrect predictions—higher KS values indicate better separation and thus more reliable confidence estimates; and (b) Expected Calibration Error (ECE guo2017calibration), which quantifies the average gap between predicted confidence and actual accuracy across binned predictions—lower ECE indicates that confidence scores accurately reflect true correctness probabilities. Together, these metrics assess both discriminative power (KS) and absolute calibration quality (ECE).
Protocol.
All methods are training-free. Prompts, temperature(0.0), tool access, and stopping rules (maximum 5 reflection iterations) are held constant across conditions; seeds are fixed for comparability. For batched settings, we vary batch size $N\in{4,8}$ and use a sequential batching strategy in the main experiments.
Appendix B Token Efficiency Results
Method SMS_Spam GQPA Winogrande
In Out Cost $\Delta$% In Out Cost $\Delta$% In Out Cost $\Delta$%
Actor
Reflection 105,936 334,429 \$3.61 196,763 421,168 \$4.70 115,788 228,050 \$2.57
BOT (4) 75,838 221,962 \$2.41 33.25% 176,381 324,153 \$3.68 21.71% 103,757 200,847 \$2.27 11.75%
BOT (8) 65,268 189,556 \$2.06 42.96% 146,350 305,647 \$3.42 27.24% 91,716 176,136 \$1.99 22.54%
Reflector
Reflection 429,383 531,556 \$6.39 602,293 125,173 \$2.76 378,801 245,424 \$3.40
BOT (4) 216,070 189,716 \$2.44 61.85% 370,783 56,221 \$1.49 45.99% 257,260 132,619 \$1.97 42.10%
BOT (8) 186,413 137,899 \$1.84 71.12% 317,082 37,507 \$1.17 57.65% 229,118 94,713 \$1.52 55.31%
Total
Reflection 535,319 865,985 \$10.00 799,056 546,341 \$7.46 494,589 473,474 \$5.97
BOT (4) 291,908 411,678 \$4.85 51.52% 547,164 380,374 \$5.17 30.68% 361,017 333,466 \$4.24 29.04%
BOT (8) 251,681 327,455 \$3.90 60.95% 463,432 343,154 \$4.59 38.48% 320,834 270,849 \$3.51 41.21%
Table 5: Input and output token usage and cost across different methods, decomposed into Actor, Reflection, and Total stages (GPT-4o). 1M input tokens cost $2.5$ and 1M output tokens cost 1010. $\Delta\%$ is the cost reduction, computed for BOT(4) and BOT(8) relative to Reflection for the same dataset and stage.
We provide a comprehensive breakdown of token usage and costs across different experimental configurations. To ensure fair comparison, all costs are normalized using the production-grade GPT-4o pricing scheme (input: \$2.50 per 1M tokens; output: \$10.00 per 1M tokens as of the experimental period).
Table 5 presents detailed token counts for each pipeline stage—Actor and Reflector—across three representative datasets. For each method, we report input tokens, output tokens, and total cost in USD. The Actor stage includes all reasoning and answer generation, while the Reflector stage encompasses evaluation, critique generation, and refinement decisions.
B.1 Efficiency Gains from Batch Processing
BoT achieves substantial efficiency improvements through two complementary mechanisms. First, instruction amortization: the Reflector’s evaluation rubric and reasoning guidelines are encoded once per batch rather than repeated for each query, saving $(N-1)\times T_{\text{inst}}$ tokens where $N$ is batch size and $T_{\text{inst}}$ is instruction length. Second, reduced iteration overhead: joint evaluation enables more accurate refinement decisions, reducing unnecessary Actor-Reflector loops.
As reported in Table 2, BoT-R achieves an average total cost reduction of 46.9% across the three benchmarks when using batch size 8. Savings are most pronounced on the SMS Spam dataset (61% reduction), where the homogeneous task structure enables highly effective batch-level evaluation. Even with batch size 4, BoT-R consistently reduces costs by 30-50% while maintaining or improving accuracy.
B.2 Stage-Level Analysis
The efficiency gains distribute differently across pipeline stages:
Actor Stage: BoT introduces minimal overhead at the Actor level, as answer generation remains largely independent. The modest savings arise from reduced refinement iterations due to more accurate Reflector feedback.
Reflector Stage: BoT delivers dramatic savings (42-71% reduction) by replacing $N$ independent reflection calls with a single joint evaluation. Larger batch sizes amplify these gains: moving from $N=4$ to $N=8$ increases Reflector savings from 42% to 57% on GPQA.
Total Cost: The combined effect yields 29-61% total cost reduction depending on dataset characteristics and batch size. These results demonstrate that BoT achieves a Pareto improvement: simultaneously enhancing both task performance (Table 1) and computational efficiency, making it particularly valuable for production deployments where cost and accuracy are both critical.
Appendix C Theoretical Analysis
This section establishes formal foundations for batch-aware reasoning in LLM-based systems. We prove that joint processing of related queries provides information-theoretic advantages over independent processing, characterize conditions under which these benefits manifest, and derive efficiency guarantees for batch-level computation.
C.1 Preliminaries and Problem Formulation
Definition C.1 (Batch reasoning problem).
Let $\mathcal{D}$ be a distribution over query-answer pairs $\mathcal{X}\times\mathcal{Y}$. A batch $B=\{(x_{i},y_{i}^{*})\}_{i=1}^{N}$ consists of $N$ instances drawn from $\mathcal{D}$. An Actor agent $\mathcal{A}$ produces initial predictions $\hat{y}_{i}^{(0)}=\mathcal{A}(x_{i})$ with reasoning traces $\rho_{i}$. The batch context is
$\mathcal{C}^{\mathrm{BoT}}=\{(x_{j},\hat{y}_{j}^{(0)},\rho_{j})\}_{j=1}^{N}.$ (6)
A Reflector agent $\mathcal{R}$ performs joint analysis over $\mathcal{C}^{\mathrm{BoT}}$ to produce for each instance $i\in[N]$:
$(\mathbf{r}_{i},u_{i},c_{i})=\mathcal{R}(\mathcal{C}^{\mathrm{BoT}},i),$ (7)
where $\mathbf{r}_{i}\in\{0,1\}$ indicates re-evaluation necessity, $u_{i}\in[0,1]$ quantifies confidence in correctness, and $c_{i}$ provides actionable critique.
Assumption C.2 (Batch coherence).
The batch exhibits structural coherence with the following properties:
  1. (a)
    Exchangeability: The joint distribution of $\{(x_{i},y_{i}^{*})\}_{i=1}^{N}$ is invariant under permutations.
  2. (b)
    Similarity structure: There exists a similarity function $\mathrm{sim}:\mathcal{X}\times\mathcal{X}\to[0,1]$ such that the average pairwise similarity satisfies $\mathbb{E}[\mathrm{sim}(x_{i},x_{j})]\geq\kappa$ for some coherence parameter $\kappa\in(0,1]$.
  3. (c)
    Error correlation: Define error indicators $e_{i}=\mathbf{1}[\hat{y}_{i}^{(0)}\neq y_{i}^{*}]$. Under coherence, errors exhibit positive correlation: $\mathrm{Cor}(e_{i},e_{j})=\rho_{e}(\kappa)>0$ for $i\neq j$, where $\rho_{e}(\cdot)$ is non-decreasing in coherence strength.
C.2 Information-Theoretic Foundation for Calibration Improvement
We first establish that batch-level processing provides strictly more information for confidence estimation than independent processing.
Theorem C.3 (Batch processing improves proper scoring rules).
Let $\mathcal{G}_{0}=\sigma(\hat{y}_{i}^{(0)},\rho_{i})$ represent the $\sigma$-algebra generated by instance $i$ alone, and $\mathcal{G}_{1}=\sigma(\hat{y}_{i}^{(0)},\rho_{i},\phi)$ where $\phi=\phi(\mathcal{C}^{\mathrm{BoT}})$ denotes batch-level statistics. For any strictly proper scoring rule $S:[0,1]\times\{0,1\}\to\mathbb{R}$ (e.g., Brier score, log-loss), the batch-aware confidence predictor
$u_{i}^{\mathrm{BoT}}=\mathbb{P}(\hat{y}_{i}^{(0)}=y_{i}^{*}\mid\mathcal{G}_{1})$ (8)
satisfies
$\mathbb{E}[S(u^{\mathrm{BoT}},z)]\leq\mathbb{E}[S(u^{\mathrm{ind}},z)],$ (9)
where $u^{\mathrm{ind}}=\mathbb{P}(\hat{y}_{i}^{(0)}=y_{i}^{*}\mid\mathcal{G}_{0})$ and $z=\mathbf{1}[\hat{y}_{i}^{(0)}=y_{i}^{*}]$. The inequality is strict when $\phi$ is informative: $I(z;\phi\mid\mathcal{G}_{0})>0$.
Proof.
By definition of conditional expectation and the tower property,
$u_{i}^{\mathrm{BoT}}$ $=\mathbb{E}[z\mid\mathcal{G}_{1}]=\mathbb{E}[\mathbb{E}[z\mid\mathcal{G}_{0}]\mid\mathcal{G}_{1}]=\mathbb{E}[u^{\mathrm{ind}}\mid\mathcal{G}_{1}].$ (10)
Since $S$ is strictly proper, the predictor $u^{\mathrm{BoT}}$ is optimal among all $\mathcal{G}_{1}$-measurable predictors. By the law of total expectation,
$\mathbb{E}[S(u^{\mathrm{BoT}},z)]$ $=\mathbb{E}[\mathbb{E}[S(u^{\mathrm{BoT}},z)\mid\mathcal{G}_{1}]]$ (11)
$\leq\mathbb{E}[\mathbb{E}[S(u^{\mathrm{ind}},z)\mid\mathcal{G}_{1}]]=\mathbb{E}[S(u^{\mathrm{ind}},z)],$ (12)
where the inequality follows from optimality of $u^{\mathrm{BoT}}$ with respect to $\mathcal{G}_{1}$. Strict inequality holds when $u^{\mathrm{BoT}}\neq u^{\mathrm{ind}}$ with positive probability, which occurs precisely when $I(z;\phi\mid\mathcal{G}_{0})>0$. ∎
Remark C.4 (Connection to Expected Calibration Error).
While Expected Calibration Error (ECE) is not a proper scoring rule, empirical calibration typically improves when confidence predictors condition on additional informative statistics. Theorem C.3 provides theoretical justification for observed ECE reductions under batch processing: by extracting cross-instance statistics $\phi$ through comparative analysis, the Reflector produces better-calibrated confidence estimates than those based solely on single-instance features.
C.3 Collective Signal Amplification for Error Detection
We now quantify how batch-level aggregation amplifies signals for error detection, explaining improved separation in confidence distributions between correct and incorrect predictions.
Proposition C.5 (Effective sample size under correlation).
Let $z_{i}=\mathbf{1}[\hat{y}_{i}^{(0)}=y_{i}^{*}]$ denote correctness indicators with $\mathbb{E}[z_{i}]=p$ and equicorrelation structure $\mathrm{Cor}(z_{i},z_{j})=\rho_{c}\in[0,1)$ for all $i\neq j$. Define the effective sample size
$N_{\mathrm{eff}}=\frac{N}{1+(N-1)\rho_{c}}.$ (13)
Then the batch-average correctness $M_{N}=\frac{1}{N}\sum_{i=1}^{N}z_{i}$ has variance
$\mathrm{Var}(M_{N})=\frac{p(1-p)}{N_{\mathrm{eff}}}.$ (14)
Furthermore:
  1. (i)
    If $\rho_{c}=O(1/N)$, then $N_{\mathrm{eff}}=\Theta(N)$ and $M_{N}$ concentrates at rate $O(1/\sqrt{N})$.
  2. (ii)
    If $\rho_{c}=\rho_{0}>0$ is constant, then $N_{\mathrm{eff}}\to 1/\rho_{0}$ as $N\to\infty$, and concentration gains saturate.
Proof.
For exchangeable binary random variables with equicorrelation $\rho_{c}$,
$\mathrm{Var}(M_{N})$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}(z_{i})+\frac{1}{N^{2}}\sum_{i\neq j}\mathrm{Cov}(z_{i},z_{j})$ (15)
$=\frac{1}{N^{2}}\cdot N\cdot p(1-p)+\frac{1}{N^{2}}\cdot N(N-1)\cdot\rho_{c}p(1-p)$ (16)
$=\frac{p(1-p)}{N}\left[1+(N-1)\rho_{c}\right]=\frac{p(1-p)}{N_{\mathrm{eff}}}.$ (17)
The asymptotic regimes follow directly from the definition of $N_{\mathrm{eff}}$. ∎
Corollary C.6 (Confidence separation for error detection).
When $N_{\mathrm{eff}}$ is large, batch-level consensus $M_{N}$ provides a reliable collective signal. Instances whose predictions deviate from consensus receive adjusted confidence scores. For fixed instance-level accuracy $p>1/2$, the Kolmogorov-Smirnov (KS) statistic measuring separation between confidence distributions of correct and incorrect predictions increases with $N_{\mathrm{eff}}$.
Remark C.7 (Optimal batch composition).
Proposition C.5 reveals a fundamental trade-off: high coherence $\kappa$ enables effective cross-instance learning but may increase error correlation $\rho_{e}$, reducing $N_{\mathrm{eff}}$. Optimal batches satisfy:
  • Sufficient similarity: $\kappa>\kappa_{\min}$ to enable pattern extraction and knowledge transfer.
  • Sufficient diversity: $\rho_{e}<0.5$ to maintain $N_{\mathrm{eff}}>0.67N$, ensuring reliable collective signals.
For batch sizes $N\in\{4,8\}$ used in practice, this yields $N_{\mathrm{eff}}\in[2.7,5.3]$ when $\rho_{e}\approx 0.3$, providing meaningful collective signal while preserving diversity.
C.4 Computational Efficiency Through Amortization
Proposition C.8 (Sublinear cost scaling).
Let $T_{\mathrm{inst}}$, $T_{\mathrm{ctx}}$, and $T_{\mathrm{out}}$ denote token counts for reflection instructions, per-instance context, and per-instance output, respectively. Independent reflection incurs total cost
$C_{\mathrm{ind}}=N\cdot(T_{\mathrm{inst}}+T_{\mathrm{ctx}}+T_{\mathrm{out}}).$ (18)
Batch-aware reflection with shared comparative analysis costs
$C_{\mathrm{BoT}}=T_{\mathrm{inst}}+N\cdot T_{\mathrm{ctx}}+S(N),$ (19)
where $S(N)$ is the joint Reflector output length. When critiques reference shared reasoning structures and cross-instance insights are reused, $S(N)$ exhibits sublinear growth: $S(N)=O(N^{\beta})$ with $\beta<1$.
C.5 Characterization of Favorable Conditions
We synthesize the preceding results to characterize when batch-aware reasoning provides advantages.
Theorem C.9 (Conditions for BoT effectiveness).
Batch-aware reasoning via BoT provides improvements over independent processing in calibration (lower ECE), error detection (higher KS statistic), and efficiency when the following conditions hold:
  1. (i)
    Coherence: Batch exhibits sufficient similarity structure with $\kappa>\kappa_{\min}$, enabling pattern extraction and knowledge transfer.
  2. (ii)
    Moderate correlation: Error correlation satisfies $\rho_{e}\in(0,0.5)$, ensuring $N_{\mathrm{eff}}>0.5N$ for collective signal reliability while preserving diversity.
  3. (iii)
    Informative batch statistics: Cross-instance features $\phi(\mathcal{C}^{\mathrm{BoT}})$ satisfy $I(z_{i};\phi\mid\mathcal{G}_{0})>0$, providing additional information beyond single-instance features.
  4. (iv)
    Adequate batch size: $N\geq N_{\min}$ for reliable collective signal extraction. For typical correlation $\rho_{e}\approx 0.3$, batch sizes $N\in\{4,8\}$ yield $N_{\mathrm{eff}}\in[2.7,5.3]$.
Under these conditions, the following guarantees hold:
  • Calibration: By Theorem C.3, batch-aware confidence $u^{\mathrm{BoT}}$ achieves lower expected loss for proper scoring rules.
  • Error detection: By Corollary C.6, confidence distributions exhibit increased separation with $N_{\mathrm{eff}}$.
  • Efficiency: By Proposition C.8, sublinear output scaling yields $C_{\mathrm{BoT}}/C_{\mathrm{ind}}<1$ for $N\geq 2$.
Remark C.10 (Failure modes and graceful degradation).
BoT degrades toward independent processing when:
  • No coherence ($\kappa\approx 0$): Instances lack shared structure; cross-instance statistics $\phi$ are uninformative.
  • High correlation ($\rho_{e}\to 1$): All instances make identical errors; $N_{\mathrm{eff}}\to 1$, eliminating collective signal benefits.
  • Insufficient size ($N<N_{\min}$): Collective signals are unreliable due to high sampling variance.
Importantly, performance degrades gracefully as $N_{\mathrm{eff}}$ decreases continuously with $\rho_{e}$, rather than exhibiting catastrophic failure.
C.6 Summary
Our theoretical analysis establishes rigorous foundations for batch-aware reasoning:
  • Information gain (Theorem C.3): Batch statistics $\phi$ provide additional information, improving calibration through optimal conditioning on $\mathcal{G}_{1}\supset\mathcal{G}_{0}$.
  • Effective sample size (Proposition C.5): Quantifies collective signal strength via $N_{\mathrm{eff}}$, explaining KS statistic improvements under moderate correlation.
  • Computational efficiency (Proposition C.8): Sublinear output scaling yields provable cost reductions for $N\geq 2$.
  • Effectiveness conditions (Theorem C.9): Characterizes when BoT succeeds, providing actionable guidance for batch construction and domain selection.
These results not only explain empirical findings but also provide principled guidelines for applying batch-aware reasoning to new domains and tasks.
Appendix D Related Work
D.1 Confidence Calibration in Large Language Models
Reliable uncertainty quantification remains critical for deploying LLMs in high-stakes applications. Modern LLMs frequently exhibit poor calibration, assigning high confidence to incorrect predictions (guo2017calibration; xiong2023uncertainty; kadavath2022language). This miscalibration persists even in state-of-the-art models (openai2023gpt4; anthropic2024claude), undermining trust in automated decision-making systems.
Existing calibration approaches fall into three categories. Post-hoc calibration methods apply temperature scaling (guo2017calibration) or Platt scaling to model outputs, but require held-out calibration sets and fail to capture semantic uncertainty (kuhn2023semantic). Sampling-based methods estimate uncertainty through self-consistency (wang2023selfconsistency), semantic entropy (kuhn2023semantic), or ensemble disagreement (chen2023reconcile). While effective, these approaches incur substantial computational overhead—self-consistency requires 10-40 samples per query—and process each instance independently, missing opportunities for cross-instance calibration. Verbalized confidence approaches directly prompt models for numerical (xiong2023uncertainty; lin2022teaching) or categorical (tian2023justsay) confidence estimates. These methods are efficient but highly sensitive to prompt formatting (si2024prompting) and often produce overconfident predictions (kadavath2022language). Recent efforts employ chain-of-thought reasoning for confidence elicitation (xiong2023uncertainty) or fine-tune models on calibration data (lin2022teaching), yet these remain per-instance techniques that cannot leverage distributional signals.
Our work introduces comparative calibration through batch processing: confidence scores are grounded in cross-instance statistics rather than isolated assessments. This approach combines the efficiency of verbalized confidence (no additional sampling) with the distributional awareness of ensemble methods, achieving superior calibration without multiplicative computational costs.
D.2 Multi-Agent Reasoning Systems
Recent work has explored sophisticated communication protocols (wu2023autogen), dynamic role allocation (hong2023metagpt), and multi-agent collaboration on complex tasks (li2023camel; qian2024chatdev). However, a fundamental limitation persists: existing multi-agent systems process queries independently. Even when multiple agents collaborate on a single query, the framework treats each query in isolation, discarding cross-instance signals. AutoGen (wu2023autogen) and MetaGPT (hong2023metagpt) enable multi-agent workflows but apply them instance-by-instance. CAMEL (li2023camel) studies role-playing conversations yet maintains per-query boundaries.
The closest work to ours is batch prompting (cheng2023batch), which groups multiple queries into a single API call for efficiency. However, batch prompting lacks reflective evaluation mechanisms and does not perform comparative analysis—it simply concatenates queries without leveraging cross-instance reasoning. Our work fundamentally differs by introducing batch-aware reflection: the Reflector explicitly performs comparative evaluation, consistency checking, and knowledge propagation across the batch.
D.3 Cross-Instance Learning
In deep learning, batch normalization (ioffe2015batch) leverages mini-batch statistics during training, while recent work explores cross-example attention (lee2019set) and set-based reasoning (zaheer2017deep). Meta-learning approaches (finn2017maml; snell2017prototypical) learn from task distributions rather than individual instances, demonstrating benefits of comparative learning.
For LLM inference, in-context learning (brown2020language) uses examples to guide reasoning, and analogical prompting (yasunaga2024large) retrieves similar cases to aid problem-solving. However, these methods rely on predefined examples or retrieval systems rather than jointly reasoning over a batch of target queries. Recent work on self-consistency with rationalization (mitchell2023enhancing) aggregates multiple reasoning paths for a single query but does not transfer knowledge across distinct queries.
BoT differs by enabling mutual information gain across queries at inference time: each query in the batch provides signal for evaluating others through comparative reflection. This creates a feedback loop where batch-level patterns inform individual assessments, analogous to how James-Stein estimation improves individual predictions through the group mean, but applied dynamically to LLM reasoning rather than static parameter estimation.
D.4 Positioning of Our Work
Our contributions address gaps in existing literature along three dimensions:
(1) Efficiency-calibration trade-off: We achieve better calibration than verbalized confidence and comparable accuracy to self-consistency while reducing costs by 46.9% (vs. per-instance reflection) rather than increasing costs 10-40× (self-consistency overhead).
(2) Cross-instance reasoning: We introduce the first multi-agent framework that explicitly performs comparative evaluation across queries, going beyond batch prompting’s simple concatenation to enable consistency checking, knowledge propagation, and distributional calibration.
(3) Training-free generality: Unlike calibration methods requiring fine-tuning (lin2022teaching) or specialized architectures, BoT is model-agnostic and integrates with existing multi-agent frameworks (Reflection, Plan-and-Act, Debate) without additional training.
Appendix E Batching Strategy Analysis
We investigate how batch composition and size influence BoT’s performance through systematic experiments across six benchmarks using GPT-4o.
E.1 Batch Size Effects
Batch FraudDet GPQA Winogrande MedQA PubMedQA SMS Spam
1(Reflection) 0.693 0.459 0.879 0.901 0.667 0.854
4 0.740 0.488 0.888 0.895 0.683 0.881
8 0.732 0.471 0.890 0.904 0.698 0.887
Table 6: Batch Size Influence on Accuracy
Table 6 presents accuracy across batch sizes $N\in\{1,4,8\}$, where $N=1$ corresponds to standard per-instance reflection. The results reveal non-monotonic relationships between batch size and performance, with optimal configurations varying by task characteristics.
Most datasets exhibit peak performance at moderate batch sizes, suggesting that batch size interacts with task structure in complex ways. Our theoretical analysis (Appendix C) predicts that larger batches increase mutual information gain by providing richer comparative context. However, three factors constrain this relationship in practice:
(1) Context window saturation. As batch size approaches model context limits, the system must compress individual rationales to fit all items. Near capacity, models produce overly concise responses that sacrifice reasoning depth for brevity, diminishing the comparative analysis benefits. For GPQA—which requires detailed scientific reasoning—this compression effect becomes apparent at $N=8$, where individual responses average 30% shorter than at $N=4$.
(2) Batch heterogeneity. When queries within a batch are too dissimilar, cross-instance signals become noisy rather than informative. Sequential batching—our default strategy—groups adjacent queries without explicit similarity filtering. For datasets with high within-domain variance (e.g., GPQA spanning biology, physics, and chemistry), larger batches increase the likelihood of mixing incompatible problem types, diluting useful comparative signals.
E.2 Semantic vs. Sequential Batching
We further study how batch composition influences BoT-R by comparing three batching strategies across six datasets. No-batch applies reflection independently to each query without any cross-instance context. Sequential batching groups queries in their original dataset order—without explicitly enforcing semantic similarity—and is used as the default strategy in our main experiments. Semantic batching clusters queries by embedding similarity using K-means over E5-Mistral-7B embeddings (wang2023improving), and forms fixed-size batches from cluster members sorted by proximity to the cluster centroid to maximize within-batch coherence. Results are summarized in Table 7.
Method FraudDet MedQA PubMedQA Winogrande SMS Spam GPQA
No-batch 0.693 0.901 0.667 0.879 0.854 0.459
Sequential 0.740 0.904 0.698 0.890 0.887 0.488
Semantic 0.768 0.902 0.697 0.897 0.902 0.486
Table 7: Accuracy comparison across batching strategies.
Robust gains from simple batching.
A key observation is that BoT-R delivers substantial improvements even under simple sequential batching. Compared to the no-batch baseline, sequential grouping improves performance on all six datasets, yielding an average relative gain of +3.87%. This demonstrates that BoT-R is not overly sensitive to imperfect batch coherence and can reliably extract useful cross-instance signals even when batches are formed without explicit semantic optimization.
Additional benefits from semantic coherence.
Semantic batching provides further gains on several datasets, particularly those where cross-instance comparison and distributional cues are informative. On FraudDet, accuracy improves from 0.740 to 0.768, and on SMS Spam from 0.887 to 0.902 when moving from sequential to semantic batching. Winogrande shows a similar trend. Averaged across all six datasets, semantic batching achieves a relative improvement of +4.83% over the no-batch baseline, exceeding that of sequential batching. These results align with our theoretical analysis: increasing within-batch coherence strengthens the informativeness of batch-level statistics, improving collective error detection and refinement decisions.
When batching strategy matters less.
For datasets characterized by shared domain knowledge or homogeneous reasoning styles, such as MedQA and PubMedQA, the difference between sequential and semantic batching is minimal. This suggests that when queries already originate from a narrow latent distribution, even naive batching satisfies the coherence conditions required for effective batch-aware reasoning, consistent with the robustness guarantees discussed in Section C.
Practical considerations.
While semantic clustering is beneficial in offline or high-throughput evaluation settings, it introduces practical constraints in streaming scenarios (e.g., online fraud detection), where queries arrive sequentially and delaying processing to form semantically coherent batches may increase latency. The strong performance of sequential batching indicates that BoT-R remains effective under such constraints. Designing adaptive cohorting strategies that balance coherence, latency, and throughput is an important direction for future work.
Appendix F MMLU Detailed Results
Table 4 reveals a systematic pattern in how BoT’s effectiveness varies across subject domains within the MMLU benchmark. We identify two distinct task categories with markedly different responses to batch-level reasoning.
Subjective and interpretive domains.
BoT-R achieves its strongest gains on humanities (+1.4% over Reflection), social sciences (+0.7%), and medicine & biology (+0.5%). These domains share three key characteristics: (1) questions often admit multiple defensible reasoning paths, (2) answer quality depends on contextual interpretation rather than strict logical derivation, and (3) comparative evaluation helps identify robust reasoning patterns across similar cases. For instance, in social science questions about policy implications or historical interpretation, batch-level reflection enables the model to distinguish well-grounded arguments from superficially plausible but contextually inconsistent reasoning.
Formal and symbolic domains.
In contrast, mathematics and physical sciences show qualitatively different behavior. While Reflection substantially improves over ReAct in these domains (+10.2% and +4.6% respectively), BoT-R exhibits marginal decline relative to Reflection (-1.2% and -1.2%). This pattern suggests that batch-level consensus can occasionally mislead reasoning in domains where correctness is determined by precise symbolic manipulation rather than comparative plausibility. In mathematical problem-solving, an incorrect but superficially consistent approach across multiple batch items may receive spurious validation through cross-instance agreement, whereas per-instance reflection focuses more directly on logical rigor.
Implications for batch composition.
These findings indicate that BoT’s effectiveness depends critically on task structure. Domains requiring interpretive judgment and context-dependent reasoning benefit from distributional signals and comparative calibration. Conversely, domains demanding exact symbolic computation may require alternative batch strategies—such as explicitly instructing the Reflector to prioritize logical correctness over cross-instance consensus, or segregating formal reasoning tasks into separate batches. Future work should investigate adaptive reflection strategies that modulate the weight given to batch-level signals based on detected task characteristics.
Subject ReAct Reflection BoT-R
Biology
anatomy 0.907 0.918 0.910
clinical_knowledge 0.913 0.898 0.913
college_biology 0.941 0.931 0.965
college_medicine 0.852 0.855 0.866
high_school_biology 0.959 0.961 0.964
human_aging 0.821 0.820 0.833
human_sexuality 0.917 0.923 0.923
medical_genetics 0.970 0.970 0.970
nutrition 0.902 0.911 0.902
professional_medicine 0.953 0.963 0.959
virology 0.573 0.564 0.582
General
global_facts 0.667 0.667 0.697
high_school_european_history 0.896 0.878 0.890
high_school_geography 0.935 0.944 0.944
high_school_us_history 0.942 0.946 0.946
high_school_world_history 0.951 0.941 0.945
miscellaneous 0.961 0.960 0.964
prehistory 0.957 0.943 0.963
Humanities
management 0.882 0.892 0.902
marketing 0.936 0.927 0.936
moral_disputes 0.871 0.868 0.881
moral_scenarios 0.707 0.767 0.767
philosophy 0.890 0.884 0.900
public_relations 0.739 0.752 0.743
Law
business_ethics 0.831 0.838 0.838
high_school_government_and_politics 0.982 0.984 0.979
international_law 0.898 0.900 0.908
jurisprudence 0.907 0.916 0.916
professional_law 0.758 0.761 0.763
us_foreign_policy 0.939 0.939 0.960
Math
abstract_algebra 0.561 0.727 0.697
college_mathematics 0.470 0.697 0.636
econometrics 0.716 0.779 0.743
elementary_mathematics 0.767 0.936 0.936
formal_logic 0.682 0.728 0.752
high_school_macroeconomics 0.906 0.915 0.913
high_school_mathematics 0.478 0.814 0.758
high_school_microeconomics 0.956 0.966 0.966
high_school_statistics 0.797 0.860 0.874
logical_fallacies 0.881 0.883 0.883
professional_accounting 0.761 0.886 0.886
Science
astronomy 0.932 0.927 0.934
college_chemistry 0.514 0.616 0.626
college_computer_science 0.767 0.869 0.859
college_physics 0.635 0.842 0.802
computer_security 0.851 0.848 0.848
conceptual_physics 0.892 0.902 0.917
electrical_engineering 0.821 0.819 0.847
high_school_chemistry 0.775 0.871 0.861
high_school_computer_science 0.934 0.960 0.970
high_school_physics 0.737 0.867 0.827
machine_learning 0.766 0.802 0.829
security_studies 0.809 0.820 0.824
Social Science
high_school_psychology 0.953 0.945 0.956
professional_psychology 0.898 0.895 0.904
sociology 0.924 0.925 0.932
world_religions 0.904 0.918 0.906
Table 8: MMLU per-subject accuracy summary grouped by category.
Appendix G Dataset: Fraud Detection
We introduce a seller-level dataset for fraud-seller detection tailored to evaluating LLM-based agent frameworks. Each instance corresponds to a single online seller and is annotated by domain experts as fraudulent (1) or non-fraudulent (0). The release contains 1,793 labeled sellers: 1,055 positives (58.8%) and 738 negatives (41.2%).
For each seller, we provide both a seller profile and one representative product profile from the seller. Seller profiles include shop_name, company_name, email_domain, and product_categories.
Product profiles include product_name, product_description, detailed_subcategory, detailed_category, and minimum_list_price_in_USD and maximum_list_price_in_USD. These fields enable models to reason over heterogeneous attributes rather than relying on free text alone.
The target label (is_fraudulent_shop) $\in\{0,1\}$ was assigned by domain experts following internal guidelines that emphasize deceptive practices and policy-violating behavior. While positively labeled cases reflect a consensus judgment of fraud, borderline cases may retain residual ambiguity typical of human annotation.
The corpus captures only the information present in the provided schema. External signals such as reputation scores, user reviews, temporal activity traces, or platform enforcement outcomes are not included. As a result, models evaluated on this dataset reason over supplied profile attributes rather than broader ecosystem signals.
Appendix H Prompts
This is the system prompt for the Fraud Detection Dataset:
You are a risk analyst expert working for an e-commerce company. Your job is to protect the platform and its customers by identifying fraudulent sellers. A fraudulent seller might engage in fraudulent activities, sell counterfeit goods, misrepresent products, or provide poor customer service. Your task is to conduct a holistic assessment based on the seller’s profile and the sample product of the seller. You are provided with the seller’s shop name, company name (some sellers may not have) and email domain, enclosed in triple backticks: - shop name: “‘SHOP_NAME“‘ - company name: “‘COMPANY_NAME“‘ - email domain: “‘EMAIL_DOMAIN“‘ You are also given the categories of products sold by the seller, enclosed in triple backticks: - product categories: “‘PRODUCT_CATEGORY“‘ You are also given the sample product of the seller, enclosed in triple backticks: - product_name: “‘PRODUCT_NAME“‘ - product_description: “‘PRODUCT_DESCRIPTION“‘ - detailed_subcategory: “‘DETAILED_SUBCATEGORY“‘ - detailed_category: “‘DETAILED_CATEGORY“‘ - min_list_price_usd: “‘MIN_LIST_PRICE_USD“‘ - max_list_price_usd: “‘MAX_LIST_PRICE_USD“‘ Note: you can use provided tools many times until you think the collected information is sufficient to answer the questions, but do avoid unnecessary tool calls. Based on the information provided and collected by tools, answer the following questions: 1. **Shop/Company Name Verification:** Based on the shop name and company name, does this appear to be a reliable/established seller? - If names seem generic, suspicious, or unfamiliar, search for the company/shop name to verify legitimacy - Note: Only use search results if they are clearly relevant to the specific shop or company name 2. **Email Domain Assessment:** Based on the email domain, does this suggest a professional business? - If using unfamiliar business domains, consider searching to check if it belongs to an established company - Note: Only use search results if they are clearly relevant to the email domain 3. **Product Information Check:** Based on the sample product name, description and the product categories, do you think it is reasonable for the seller to sell the products in the shop? 4. **Product Price Verification:** Does the product pricing seem reasonable for the category? - If pricing appears suspiciously low or high, search for typical market prices of similar products 5. Based on all the information, do you think this seller is a fraudulent seller? Assign a confidence score: rate your confidence in the assessment. Return your response in a single JSON object with the following keys: - ‘is_fraudulent_shop‘: (boolean) ‘true‘ if the shop exhibits indicators of fraudulent operations, otherwise ‘false‘. - ‘confidence_score‘: (float) A score from 0.0 to 1.0 indicating your confidence in the assessment. - ‘summary_reasoning‘: (string) Comprehensive explanation of your fraud assessment, including all factors that led to your conclusion.
This is the system prompt for the reflector:
You are a reflection agent to help refine the answers. Here are ¡¡N¿¿ questions, each with the previous model’s answer. For each, critique the model answer for accuracy, completeness, and reasoning, comparing across all answers and their reasoning paths in the batch to identify areas for improvement and give a peer confidence score to quantify how possible the answer is correct. Make sure you understand each question-answer pair and give detailed explanations to them, Carefully decide if a reevaluation is needed for each case. For each, provide: (1) whether to trigger reevaluation (true/false) and improve answer, (2) summary assessment, (3) peer confidence score for the current answer(0.0-1.0), (4) suggestions for improvement(empty if reevaluation is false). Output a JSON list, one entry per question, strictly in format: “response:{trigger_reevaluation: bool, summary_comment: str, confidence_score: float(0.0-1.0), suggestions: str}]”

### v2

@@
 Our theoretical and experimental analysis reveals when and why batch-aware reasoning benefits LLM systems.
+Our code is available at https://github.com/xuanyang19/BoT.
 Batch-of-Thought: Cross-Instance Learning for Enhanced LLM Reasoning
@@
 1 Introduction
-Large Language Models (LLMs) openai2023gpt4; madaan2023selfrefine; wei2022cot; shinn2023reflexion; yao2023react have achieved strong performance across diverse tasks and are increasingly applied in domains such as medical reasoning, question answering, and scientific problem solving (Singhal2025TowardEMA; mcduff2025towards; nori2025sequential; haas2025simpleqa; wang2023scibench; sun2024scieval).
-However, producing reliable answers with well-calibrated confidence remains a challenge  xiong2023uncertainty; ji2023hallucination_survey.
+Large Language Models (LLMs) OpenAI (2023); Madaan et al.
+(2023); Wei et al.
+(2022); Shinn et al.
+(2023); Yao et al.
+(2023) have achieved strong performance across diverse tasks and are increasingly applied in domains such as medical reasoning, question answering, and scientific problem solving (Singhal et al., 2025; McDuff et al., 2025; Nori et al., 2025; Haas et al., 2025; Wang et al., 2023b; Sun et al., 2024).
+However, producing reliable answers with well-calibrated confidence remains a challenge  Xiong et al.
+(2023); Ji et al.
+(2023).
 LLMs often assign high confidence to incorrect answers, which undermines their practical deployment in high-stakes applications where accuracy and reliable uncertainty quantification are essential.
@@
 A Reflector then jointly evaluates all responses through comparative analysis, determining whether each should be finalized or refined with feedbacks.
-Multi-agent LLM systems (li2023camel; chan2023chateval; guo2024large) extend single-model capabilities through specialized roles and iterative refinement.
+Multi-agent LLM systems (Li et al., 2023; Chan et al., 2023; Guo et al., 2024) extend single-model capabilities through specialized roles and iterative refinement.
 Despite architectural diversity, existing approaches share a fundamental limitation: they process queries independently.
@@
 Examining a cohort simultaneously reveals recurring suspicious patterns, distinguishes legitimate domain practices from genuine anomalies, and enables comparative evidence assessment.
-This principle mirrors James–Stein estimation james1961quadratic; stein1956inadmissibility; efron1977steinsparadox: pooling information across similar instances improves individual estimates through shrinkage toward the cohort distribution (see Appendix C for theoretical analysis).
-We instantiate BoT within a multi-agent reflection architecture madaan2023selfrefine; shinn2023reflexion, termed BoT-R, though the principle generalizes to other frameworks including Plan-and-Act Erdogan2025PlanAndAct and Multi-Agent Debate Du2023MultiAgentDebate; Liang2023MAD.
+This principle mirrors James–Stein estimation James and Stein (1961); Stein (1956); Efron and Morris (1977): pooling information across similar instances improves individual estimates through shrinkage toward the cohort distribution (see Appendix C for theoretical analysis).
+We instantiate BoT within a multi-agent reflection architecture Madaan et al.
+(2023); Shinn et al.
+(2023), termed BoT-R, though the principle generalizes to other frameworks including Plan-and-Act Erdogan et al.
+(2025) and Multi-Agent Debate Du et al.
+(2023); Liang et al.
+(2023).
 In BoT-R, an Actor generates answer-rationale pairs for a batch of queries, then a Reflector performs joint evaluation through comparative analysis—identifying inconsistencies, extracting shared domain knowledge, assessing relative quality, and suggesting refinements.
@@
 Actor $\mathcal{A}$.
-A ReAct agent yao2023react that interleaves reasoning traces with tool execution to generate answer-rationale pairs.
+A ReAct agent Yao et al.
+(2023) that interleaves reasoning traces with tool execution to generate answer-rationale pairs.
 At iteration $t$, given query $x_{i}$ and optional critique $c_{i}^{(t-1)}$ from the previous round:
@@
 Reflector $\mathcal{R}$.
-A reflection agent madaan2023selfrefine; shinn2023reflexion that evaluates context $\mathcal{C}^{(t)}$ containing all current answer-rationale pairs. For each query $i$, it produces:
+A reflection agent Madaan et al. (2023); Shinn et al. (2023) that evaluates context $\mathcal{C}^{(t)}$ containing all current answer-rationale pairs. For each query $i$, it produces:
 $(r_{i}^{(t)},u_{i}^{(t)},c_{i}^{(t)})=\mathcal{R}(\mathcal{C}^{(t)},i),$ (2)
@@
 Datasets.
-Our evaluation covers diverse reasoning and decision-making tasks: GPQA (rein2024gpqa), WinoGrande–debiased (sakaguchi2021winogrande), PubMedQA (jin2019pubmedqa), MedQA (USMLE) (jin2021disease), MMLU (hendrycks2020measuring), SMS Spam Detection (yang2024ad), and a newly curated dataset fraud-seller detection dataset (Appendix G).
+Our evaluation covers diverse reasoning and decision-making tasks: GPQA (Rein et al., 2024), WinoGrande–debiased (Sakaguchi et al., 2021), PubMedQA (Jin et al., 2019), MedQA (USMLE) (Jin et al., 2021), MMLU (Hendrycks et al., 2020), SMS Spam Detection (Yang et al., 2024), and a newly curated dataset fraud-seller detection dataset (Appendix G).
 Together, these benchmarks span scientific reasoning, commonsense inference, biomedical QA, broad academic knowledge, and real-world anomaly detection.
 Metrics.
-We evaluate (i) task accuracy, (ii) token efficiency (input token count, output token count, and total), and (iii) confidence calibration using two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation between the confidence distributions of correct vs.
-incorrect predictions, and (b) Expected Calibration Error (ECE; guo2017calibration).
+We evaluate (i) task accuracy, (ii) token efficiency (input token count, output token count, and total), and (iii) confidence calibration using two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic Smirnov (1939) between the confidence distributions of correct vs.
+incorrect predictions, and (b) Expected Calibration Error (ECE; Guo et al.
+(2017)).
 Baselines.
-To isolate BoT’s contribution, we compare against two training-free reasoning baselines: ReAct (yao2023react), which performs standard single-instance reasoning with optional tool augmentation, and Reflection (shinn2023reflexion; madaan2023selfrefine), a multi-agent framework employing per-instance self-critique and revision.
+To isolate BoT’s contribution, we compare against two training-free reasoning baselines: ReAct (Yao et al., 2023), which performs standard single-instance reasoning with optional tool augmentation, and Reflection (Shinn et al., 2023; Madaan et al., 2023), a multi-agent framework employing per-instance self-critique and revision.
 We then augment the same Actor with BoT’s joint, batch-aware reflection to obtain BoT-R.
@@
 Models.
-We report results from both API and open-source models: GPT-4o-2024-11-20 hurst2024gpt, Llama-3.3-70B llama3 and Qwen3-Next-80B yang2025qwen3.
-For the Fraud Detection dataset, we enable Brave Search and the Brave Summarizer as external tools for retrieval and grounding brave_api_docs
+We report results from both API and open-source models: GPT-4o-2024-11-20 Hurst et al.
+(2024), Llama-3.3-70B Dubey et al.
+(2024) and Qwen3-Next-80B Yang et al.
+(2025).
+For the Fraud Detection dataset, we enable Brave Search and the Brave Summarizer as external tools for retrieval and grounding Brave Software, Inc.
+(2025)
 SMS_Spam GPQA Winogrande
@@
 Confidence Calibration Metrics.
-We use two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation, which measures the maximum difference between the cumulative distributions of confidence scores for correct versus incorrect predictions—higher KS values indicate better separation and thus more reliable confidence estimates; and (b) Expected Calibration Error (ECE guo2017calibration), which quantifies the average gap between predicted confidence and actual accuracy across binned predictions—lower ECE indicates that confidence scores accurately reflect true correctness probabilities.
+We use two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic Smirnov (1939), which measures the maximum difference between the cumulative distributions of confidence scores for correct versus incorrect predictions—higher KS values indicate better separation and thus more reliable confidence estimates; and (b) Expected Calibration Error (ECE Guo et al.
+(2017)), which quantifies the average gap between predicted confidence and actual accuracy across binned predictions—lower ECE indicates that confidence scores accurately reflect true correctness probabilities.
 Together, these metrics assess both discriminative power (KS) and absolute calibration quality (ECE).
@@
 Reliable uncertainty quantification remains critical for deploying LLMs in high-stakes applications.
-Modern LLMs frequently exhibit poor calibration, assigning high confidence to incorrect predictions (guo2017calibration; xiong2023uncertainty; kadavath2022language).
-This miscalibration persists even in state-of-the-art models (openai2023gpt4; anthropic2024claude), undermining trust in automated decision-making systems.
+Modern LLMs frequently exhibit poor calibration, assigning high confidence to incorrect predictions (Guo et al., 2017; Xiong et al., 2023; Kadavath et al., 2022).
+This miscalibration persists even in state-of-the-art models (OpenAI, 2023; Anthropic, 2024), undermining trust in automated decision-making systems.
 Existing calibration approaches fall into three categories.
-Post-hoc calibration methods apply temperature scaling (guo2017calibration) or Platt scaling to model outputs, but require held-out calibration sets and fail to capture semantic uncertainty (kuhn2023semantic).
-Sampling-based methods estimate uncertainty through self-consistency (wang2023selfconsistency), semantic entropy (kuhn2023semantic), or ensemble disagreement (chen2023reconcile).
+Post-hoc calibration methods apply temperature scaling (Guo et al., 2017) or Platt scaling to model outputs, but require held-out calibration sets and fail to capture semantic uncertainty (Kuhn et al., 2023).
+Sampling-based methods estimate uncertainty through self-consistency (Wang et al., 2023c), semantic entropy (Kuhn et al., 2023), or ensemble disagreement (Chen et al., 2024).
 While effective, these approaches incur substantial computational overhead—self-consistency requires 10-40 samples per query—and process each instance independently, missing opportunities for cross-instance calibration.
-Verbalized confidence approaches directly prompt models for numerical (xiong2023uncertainty; lin2022teaching) or categorical (tian2023justsay) confidence estimates.
-These methods are efficient but highly sensitive to prompt formatting (si2024prompting) and often produce overconfident predictions (kadavath2022language).
-Recent efforts employ chain-of-thought reasoning for confidence elicitation (xiong2023uncertainty) or fine-tune models on calibration data (lin2022teaching), yet these remain per-instance techniques that cannot leverage distributional signals.
+Verbalized confidence approaches directly prompt models for numerical (Xiong et al., 2023; Lin et al., 2022) or categorical (Tian et al., 2023) confidence estimates.
+These methods are efficient but highly sensitive to prompt formatting (Si et al., 2024) and often produce overconfident predictions (Kadavath et al., 2022).
+Recent efforts employ chain-of-thought reasoning for confidence elicitation (Xiong et al., 2023) or fine-tune models on calibration data (Lin et al., 2022), yet these remain per-instance techniques that cannot leverage distributional signals.
 Our work introduces comparative calibration through batch processing: confidence scores are grounded in cross-instance statistics rather than isolated assessments.
@@
 D.2 Multi-Agent Reasoning Systems
-Recent work has explored sophisticated communication protocols (wu2023autogen), dynamic role allocation (hong2023metagpt), and multi-agent collaboration on complex tasks (li2023camel; qian2024chatdev).
+Recent work has explored sophisticated communication protocols (Wu et al., 2023), dynamic role allocation (Hong et al., 2023), and multi-agent collaboration on complex tasks (Li et al., 2023; Qian et al., 2024).
 However, a fundamental limitation persists: existing multi-agent systems process queries independently.
 Even when multiple agents collaborate on a single query, the framework treats each query in isolation, discarding cross-instance signals.
-AutoGen (wu2023autogen) and MetaGPT (hong2023metagpt) enable multi-agent workflows but apply them instance-by-instance.
-CAMEL (li2023camel) studies role-playing conversations yet maintains per-query boundaries.
-The closest work to ours is batch prompting (cheng2023batch), which groups multiple queries into a single API call for efficiency.
+AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2023) enable multi-agent workflows but apply them instance-by-instance.
+CAMEL (Li et al., 2023) studies role-playing conversations yet maintains per-query boundaries.
+The closest work to ours is batch prompting (Cheng et al., 2023), which groups multiple queries into a single API call for efficiency.
 However, batch prompting lacks reflective evaluation mechanisms and does not perform comparative analysis—it simply concatenates queries without leveraging cross-instance reasoning.
@@
 D.3 Cross-Instance Learning
-In deep learning, batch normalization (ioffe2015batch) leverages mini-batch statistics during training, while recent work explores cross-example attention (lee2019set) and set-based reasoning (zaheer2017deep).
-Meta-learning approaches (finn2017maml; snell2017prototypical) learn from task distributions rather than individual instances, demonstrating benefits of comparative learning.
-For LLM inference, in-context learning (brown2020language) uses examples to guide reasoning, and analogical prompting (yasunaga2024large) retrieves similar cases to aid problem-solving.
+In deep learning, batch normalization (Ioffe and Szegedy, 2015) leverages mini-batch statistics during training, while recent work explores cross-example attention (Lee et al., 2019) and set-based reasoning (Zaheer et al., 2017).
+Meta-learning approaches (Finn et al., 2017; Snell et al., 2017) learn from task distributions rather than individual instances, demonstrating benefits of comparative learning.
+For LLM inference, in-context learning (Brown et al., 2020) uses examples to guide reasoning, and analogical prompting (Yasunaga et al., 2024) retrieves similar cases to aid problem-solving.
 However, these methods rely on predefined examples or retrieval systems rather than jointly reasoning over a batch of target queries.
-Recent work on self-consistency with rationalization (mitchell2023enhancing) aggregates multiple reasoning paths for a single query but does not transfer knowledge across distinct queries.
+Recent work on self-consistency with rationalization (Mitchell et al., 2023) aggregates multiple reasoning paths for a single query but does not transfer knowledge across distinct queries.
 BoT differs by enabling mutual information gain across queries at inference time: each query in the batch provides signal for evaluating others through comparative reflection.
@@
 (2) Cross-instance reasoning: We introduce the first multi-agent framework that explicitly performs comparative evaluation across queries, going beyond batch prompting’s simple concatenation to enable consistency checking, knowledge propagation, and distributional calibration.
-(3) Training-free generality: Unlike calibration methods requiring fine-tuning (lin2022teaching) or specialized architectures, BoT is model-agnostic and integrates with existing multi-agent frameworks (Reflection, Plan-and-Act, Debate) without additional training.
+(3) Training-free generality: Unlike calibration methods requiring fine-tuning (Lin et al., 2022) or specialized architectures, BoT is model-agnostic and integrates with existing multi-agent frameworks (Reflection, Plan-and-Act, Debate) without additional training.
 Appendix E Batching Strategy Analysis
@@
 Sequential batching groups queries in their original dataset order—without explicitly enforcing semantic similarity—and is used as the default strategy in our main experiments.
-Semantic batching clusters queries by embedding similarity using K-means over E5-Mistral-7B embeddings (wang2023improving), and forms fixed-size batches from cluster members sorted by proximity to the cluster centroid to maximize within-batch coherence.
+Semantic batching clusters queries by embedding similarity using K-means over E5-Mistral-7B embeddings (Wang et al., 2023a), and forms fixed-size batches from cluster members sorted by proximity to the cluster centroid to maximize within-batch coherence.
 Results are summarized in Table 7.